An airplane left an airport flying at at 180 mph. A jet that flies at 333 mph left 1 hour later. The jet follows tthe same route as the airplane on parallel altitudes. How many hours will it take the jet to catch up with the airplane?
Answer by nerdybill(6958)

An airplane left an airport flying at at 180 mph. A jet that flies at 333 mph left 1 hour later. The jet follows tthe same route as the airplane on parallel altitudes. How many hours will it take the jet to catch up with the airplane?
.
This problem needs you to apply the "distance formula":
d = rt
where
d is distance
r is rate or speed
t is time
.
"distance airplane traveled" = "distnce jet traveled"
.
Let x = hours "airplane" was flying
x-1 = hours "jet" was flying
.
180x = 333(x-1)
180x = 333x - 333
180x + 333 = 333x
333 = 153x
333/153 = x
2.176 hours = x
.